OpenStudy (wwe123):

Quadratic Equation------ if the roots of ax^2+x+b=0 be real and unequal ,show that the roots of (x^2+1)/x =4√ab are imaginary.

OpenStudy (wwe123):

if the roots of ax^2+x+b=0 be real and unequal ,show that the roots of (x^2+1)/x =4√ab are imaginary roots.

hartnn (hartnn):

since roots of ax^2+x+b=0 be real, we can say 1-4ab>0 got this ?

OpenStudy (wwe123):

yes

hartnn (hartnn):

now, x^2 - 4\(\sqrt{ab}\)x+1=0 has, determinant as 16ab-4 ok?

OpenStudy (wwe123):

i do not understand

hartnn (hartnn):

which part ? did u get how i got 1-4ab>0 ??

OpenStudy (wwe123):

yes

hartnn (hartnn):

how did i get it ?

OpenStudy (wwe123):

this part x^2 - 4ab√x+1=0 has, determinant as 16ab-4 ok?

hartnn (hartnn):

i meant u know how to find determinant of a quadratic equation ?

OpenStudy (wwe123):

yes i know how .......

hartnn (hartnn):

so find the determinant of 2nd equation. x^2 - 4\(\sqrt{ab}\)x+1=0

hartnn (hartnn):

or u did not get how i got that equation^ ?

hartnn (hartnn):

(x^2+1)/x =4√ab (x^2+1) =4√ab x x^2+1 -4√ab x=0

OpenStudy (wwe123):

is 4ab >-1 is this correct

hartnn (hartnn):

1-4ab>0 1>4ab 4ab<1 16ab<4 16ab-4<0 determinant is negative-->roots are imaginary.